### v3

# Taming OOD Actions for Offline Reinforcement Learning: An Advantage-Based Approach

Abstract
Offline reinforcement learning (RL) aims to learn decision-making policies from fixed datasets without online interactions, providing a practical solution where online data collection is expensive or risky. However, offline RL often suffers from distribution shift, resulting in inaccurate evaluation and substantial overestimation on out-of-distribution (OOD) actions. To address this, existing approaches incorporate conservatism by indiscriminately discouraging all OOD actions, thereby hindering the agent’s ability to generalize and exploit beneficial ones. In this paper, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically evaluates OOD actions using the batch-optimal value function. Based on this evaluation, ADAC defines an advantage function to modulate the Q-function update, enabling more precise assessment of OOD action quality. We design a custom PointMaze environment and collect datasets to visually reveal that advantage modulation can effectively identify and select superior OOD actions. Extensive experiments show that ADAC achieves state-of-the-art performance on almost all tasks in the D4RL benchmark, with particularly clear margins on the more challenging tasks. Our code is available at https://github.com/NUS-CORE/adac.
1 Introduction
Offline reinforcement learning (RL) lange2012batch ; levine2020offline focuses on learning decision-making policies solely from previously collected datasets, entirely circumventing the need for online interactions with the environment. This paradigm is particularly appealing for applications where online data collection is prohibitively expensive or poses safety concerns, thereby fueling its growing adoption kalashnikov2018scalable ; prudencio2023survey . While conceptually attractive, offline RL often suffers from distribution shift between the behavior policy and the learned policy. The policy evaluation on out-of-distribution (OOD) actions is prone to extrapolation error fujimoto2019off , which can be amplified through bootstrapping, leading to significant overestimation kumar2020conservative .
To mitigate overestimation, a common strategy in offline RL is to incorporate conservatism into algorithm design. Value-based methods kumar2020conservative ; kostrikov2021offline ; lyu2022mildly achieve this by learning a pessimistic value function that reduces the estimated value of OOD actions to discourage their selection. Alternatively, policy-based methods fujimoto2021minimalist ; fujimoto2019off ; wang2022diffusion enforce conservatism by constraining the learned policy to remain close to the behavior policy, thereby avoiding querying OOD actions. Similarly, conditional sequence modeling approaches chen2021decision ; janner2022planning ; ajay2022conditional inherently induce conservative behavior by restricting the policy to imitate the behavior contained in the offline dataset. In a distinct manner, model-based approaches kidambi2020morel ; yu2020mopo ; sun2023model ensure conservatism by learning a pessimistic dynamics model where uncertainty-based penalization systematically underestimates the value of OOD actions.
However, while conservatism is celebrated in offline RL, existing methods achieve it by indiscriminately discouraging all OOD actions, thereby hindering their capacity for generalization. Offline datasets, in practice, are usually characterized by sub-optimal trajectories and narrow state-action coverage. Consequently, an effective offline RL algorithm should possess the ability to stitch sub-optimal trajectories to generate the best possible trajectory supported by the dataset, and even extrapolate beyond the dataset to identify potentially beneficial actions. Such indiscriminate discouragement, however, severely impedes the agent’s ability to generalize and achieve high performance. This naturally leads to a fundamental question: can we develop a method that reliably distinguishes between undesirable and beneficial OOD actions, thereby balancing conservatism and generalization?
To this end, we propose Advantage-based Diffusion Actor-Critic (ADAC), a novel method that systematically assesses the quality of OOD actions, selectively encourages beneficial ones, and discourages risky ones to maintain conservatism. Our key idea builds on the insight that the optimal value function provides a natural baseline for evaluating state desirability. Specifically, we regard an OOD action as advantageous if it can move the current state to a successor state whose value, under the optimal value function, exceeds that of any reachable state under the behavior policy. Since the true optimal value function is inaccessible from offline data, we adopt the batch-optimal value function ($V_{\mu}^{\ast}$ in Figure 1) as a principled surrogate. We show theoretically that the batch-optimal value function can be reliably approximated via expectile regression on the dataset. Based on this approximation, we define an advantage function to assess the desirability of actions, which is then used to modulate the temporal difference (TD) target during Q-function (critic) learning. The policy (actor) is parameterized as a diffusion model due to its strong expressiveness ho2020denoising ; wang2022diffusion and is updated under the guidance of the learned Q-function. We summarize the full architecture of ADAC in Figure 1. To the best of our knowledge, ADAC represents a pioneering attempt to systematically evaluate OOD actions, enabling the selection of beneficial ones. As a result, ADAC achieves state-of-the-art (SOTA) performance on the majority of the D4RL fu2020d4rl benchmark tasks.
Figure 1: Full architecture of ADAC. Training involves learning a diffusion actor and a Q-critic with advantage modulation, while inference deploys the learned diffusion policy.
2 Preliminaries
Offline Reinforcement Learning. RL problems are commonly formulated within the framework of a Markov Decision Process (MDP), defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},r,\rho_{0},P,\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ represents the action space, $r(\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\to[-R_{\max},R_{\max}]$ is a bounded reward function, $\rho_{0}(\bm{s})$ specifies the initial state distribution, $P(\bm{s^{\prime}}\,|\,\bm{s},\bm{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}_{+}$ defines the transition dynamics, and $\gamma\in(0,1)$ is the discount factor sutton2018reinforcement .
A policy $\pi(\cdot|\bm{s})$ maps a given state $\bm{s}$ to a probability distribution over the action space. The value function of a state $\bm{s}$ under a policy $\pi$ is the expected cumulative return when starting in $\bm{s}$ and following $\pi$ thereafter, i.e., $V^{\pi}(\bm{s})=\mathbb{E}_{\bm{a_{t}}\sim\pi(\cdot|\bm{s_{t}})}\big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s_{t}},\bm{a_{t}})\,\big{|}\,\bm{s_{0}}=\bm{s}\big{]},$ where the expectation takes over the randomness of the policy $\pi$ and the transition function $P$. The optimal state value function satisfies the following Bellman optimality equation:
$V^{\ast}(\bm{s})=\max\limits_{\bm{a}\in\mathcal{A}}\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot\,|\,\bm{s},\bm{a})}\big{[}r(\bm{s},\bm{a})+\gamma V^{\ast}(\bm{s^{\prime}})\big{]}.$ (1)
The action-value function (Q-function) is the expected cumulative return when starting from state $\bm{s}$, taking action $\bm{a}$, and following $\pi$ thereafter: $Q^{\pi}(\bm{s},\bm{a})=\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s_{t}},\bm{a_{t}})\,\big{|}\,\bm{s_{0}}=\bm{s},\bm{a_{0}}=\bm{a}\big{]}$. The goal of RL is to learn a policy $\pi(\cdot\,|\,\bm{s})$ that maximizes the following expected cumulative long-term reward:
[…]
As shown in Eq. (2), the classical RL framework requires online interactions with the environment $P$ during training. In contrast, the offline RL learns only from a fixed dataset $\mathcal{D}=\{(\bm{s},\bm{a},r,\bm{s^{\prime}})\}$ collected by the behavior policy $\mu(\cdot\,|\,\bm{s})$, where $\bm{s}$, $\bm{a}$, $r$, and $\bm{s^{\prime}}$ denote the state, action, reward, and next state, respectively. That is, it aims to find the best possible policy solely from $\mathcal{D}$ without additional interactions with the environment.
Diffusion Model. Diffusion models sohl2015deep ; ho2020denoising ; song2020score consist of a forward process that corrupts data with noise and a reverse process that reconstructs data from noise. Specifically, the forward process is conducted by gradually adding Gaussian noise to samples $\bm{x_{0}}$ from an unknown data distribution $p_{\theta}(\bm{x_{0}})$, formulated as:
$q(\bm{x_{1:T}}\,|\,\bm{x_{0}}):=\prod_{t=1}^{T}q(\bm{x_{t}}\,|\,\bm{x_{t-1}}),\quad q(\bm{x_{t}}\,|\,\bm{x_{t-1}}):=\mathcal{N}(\bm{x_{t}};\sqrt{1-\beta_{t}}\bm{x_{t-1}},\beta_{t}\bm{I}),$ (3)
where $T$ denotes the total number of diffusion steps, and $\alpha_{t}$ controls the variance of the added noise at each step $t$. The reverse process is modeled as $p_{\theta}(\bm{x_{0:T}}):=p(\bm{x_{T}})\prod_{t=1}^{T}p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}})$, and is trained by maximizing the evidence lower bound (ELBO) jordan1999introduction ; blei2017variational : $\mathbb{E}q\left[\log\frac{p_{\theta}(\bm{x_{0:T}})}{q(\bm{x_{1:T}}\mid\bm{x_{0}})}\right]$. After training, samples can be generated by first drawing $\bm{x_{T}}\sim p(\bm{x_{T}})$ and then sequentially applying the learned reverse transitions to obtain $\bm{x_{0}}$. For conditional generation tasks, the reverse process can be extended to model $p_{\theta}(\bm{x_{t-1}}\,|\,\bm{x_{t}},c)$, where $c$ denotes the conditioning information.
Expectile Regression. Expectile regression has been extensively studied in econometrics newey1987asymmetric and recently introduced in offline RL kostrikov2021offline . The $\tau$-expectile (with $\tau\in(0,1)$) of a real-valued random variable $\bm{X}$ is defined as the solution to the asymmetric least squares problem:
$\mathop{\arg\min}\limits_{y\in\mathbb{R}}\mathbb{E}_{x\sim X}\big{[}L_{2}^{\tau}(x-y)\big{]},$ (4)
where $L_{2}^{\tau}(u)=|\tau-\mathds{1}(u<0)|u^{2}$. Therefore, $\tau=0.5$ corresponds to the standard mean squared error loss, while $\tau>0.5$ downweights the contributions of $x$ values smaller than $y$ and assigns greater weight to larger values. As $\tau\to 1$, the solution to Eq. (4) asymptotically approaches the maximum value within the support of $X$.
3 Theoretical Foundation for Advantage-Based Evaluation
In this section, we establish the theoretical foundation for our method.
3.1 Batch-Optimal Value Function Learning via Expectile Regression
In this subsection, we introduce an advantage function designed to evaluate all actions, including OOD actions. To this end, we first approximate the optimal value function, as it serves as a natural reference for assessing state desirability. Since the optimal value function satisfies the Bellman optimality equation defined in Eq. (1), we adopt expectile regression to approximate the maximum operator, while replacing the expectation with empirical samples drawn from the offline dataset. This naturally suggests the following regression problem:
$\mathcal{L}(V)=\mathbb{E}_{(\bm{s},\bm{a},r,\bm{s^{\prime}})\sim\mathcal{D}}\big{[}L_{2}^{\tau}(r(\bm{s},\bm{a})+\gamma V(\bm{s^{\prime}})-V(\bm{s}))\big{]}.$ (5)
Notably, this formulation entirely circumvents the need to query unseen actions. Let $V_{\tau}(\bm{s})$ denote the minimizer of $\mathcal{L}(V)$ defined in Eq. (5). The following proposition characterizes its analytical form.
Proposition 3.1.
The minimizer $V_{\tau}(\bm{s})$ satisfies
$V_{\tau}(\bm{s})=\mathbb{E}^{\tau}_{\bm{a}\sim\mu(\cdot|\bm{s}),\ \bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}r(\bm{s},\bm{a})+\gamma V_{\tau}(\bm{s^{\prime}})\big{]},$ (6)
where $\mathbb{E}^{\tau}_{x\sim X}[x]$ denotes the $\tau$-expectile of a real-valued random variable $X$.
We further define $V_{\mu}^{\ast}(\bm{s})$ as the batch-optimal value function, which satisfies
$V_{\mu}^{\ast}(\bm{s}):=\max\limits_{\bm{a}\sim\mu(\cdot|\bm{s})}\bigg{[}r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}V^{\ast}_{\mu}(\bm{s^{\prime}})\big{]}\bigg{]}.$ (7)
The batch-optimal value function defined in Eq. (7) differs from the optimal value function in Eq. (1) in that it restricts the maximization to actions observed in the batch dataset. Therefore, it serves as a principled approximation to the optimal value function over the support of the batch data. The next proposition establishes the relation between the minimizer $V_{\tau}(\bm{s})$ and the batch-optimal value function $V_{\mu}^{\ast}(\bm{s})$.
Proposition 3.2.
The solution $V_{\tau}(\bm{s})$ to Eq. (5) is uniformly bounded and monotonically non-decreasing with respect to $\tau$. Its pointwise limit as $\tau\to 1$, denoted by $\bar{V}(\bm{s})$, satisfies
$\bar{V}(\bm{s})=\max_{a\sim\mu(\cdot|\bm{s})}\left[r(\bm{s},\bm{a})+\gamma\max_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bar{V}(\bm{s^{\prime}})\right].$ (8)
In the case where the transition dynamics P are deterministic, the limit coincides with the batch-optimal value function, i.e.,
$\lim_{\tau\to 1}V_{\tau}(\bm{s})=\bar{V}(\bm{s})=V_{\mu}^{\ast}(\bm{s}).$
According to Proposition 3.2, under deterministic transitions, $V_{\tau}(\bm{s})$ converges to $V_{\mu}^{\ast}(\bm{s})$ as $\tau\to 1$, thereby offering a precise approximation to $V^{\ast}(\bm{s})$ when the dataset $\mathcal{D}$ provides complete coverage over the action space. In contrast, under stochastic transitions, the limit $\bar{V}(\bm{s})$ involves a maximization over next states, which can lead to $\bar{V}(\bm{s})\geq V_{\mu}^{\ast}(\bm{s})$. Nonetheless, since $V_{\tau}(\bm{s})$ is monotonically non-decreasing in $\tau$, choosing $\tau<1$ yields a practical approximation to $V_{\mu}^{\ast}(\bm{s})$. This is further corroborated by our experiments, which show the best performance occurs for $\tau$ in $(0.6,0.8)$.
The above analysis justifies the approximation of the optimal value function via Eq. (5), which is subsequently used to assess the quality of OOD actions.
3.2 Advantage-Based Operator Analysis
As shown in the previous subsection, given a specific choice of $\tau$, we denote the learned value function from Eq. (5) as $V(\bm{s})$, which approximates the batch-optimal value function. Our idea is based on the observation that the quality of an action can be assessed by whether it transitions to a more desirable next state. Accordingly, we define the advantage of an action $\bm{a}$ (possibly OOD) over the behavior policy at state $\bm{s}$ as
$A(\bm{a}|\bm{s}):=\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot\mid\bm{s},\bm{a})}V(\bm{s^{\prime}})-\mathop{\text{Quantile}}\limits_{\kappa}\left(\left\{\mathbb{E}_{\bm{s^{\prime}_{i}}\sim P(\cdot\mid\bm{s},\bm{a_{i}})}V(\bm{s^{\prime}_{i}})\right\}_{i=1}^{N}\right),\quad\bm{a_{i}}\sim\mu(\cdot|\bm{s}),$ (9)
where $\{\bm{a_{i}}\}_{i=1}^{N}$ are $N$ actions independently sampled from the behavior policy $\mu(\cdot|\bm{s})$. In our experiments, we fix $N\equiv 25$ to balance performance and computational efficiency. $\mathop{\text{Quantile}}_{\kappa}(\cdot)$ denotes the $\kappa$-th quantile of the expected next-state values induced by behavior policy actions.
Under this definition, an action $\bm{a}$ is considered advantageous if it leads to a next state with a higher expected value than the selective threshold defined by the $\kappa$-th quantile. A positive advantage indicates that the action is favored, while a negative advantage indicates that the action is penalized. The parameter $\kappa$ controls the level of conservatism: larger values lead to higher thresholds and encourage conservatism, while smaller values promote optimism by more readily rewarding unseen actions. Notably, all the components in Eq. (9) are learned solely from the offline dataset, providing a reliable basis for evaluating the quality of OOD actions.
Building on the preceding development, we now augment the standard Bellman operator using the advantage function. Specifically, we introduce the following advantage-based Bellman operator:
$\mathcal{T}_{A}^{\pi_{\theta}}Q(\bm{s},\bm{a})=r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a}),\,\bm{a^{\prime}}\sim\pi_{\theta}}\big{[}Q(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}|\bm{s^{\prime}})\big{]},$ (10)
where $\lambda$ is a scaling coefficient that modulates the influence of the advantage function.
Offline RL algorithms based on the standard Bellman backup suffer from action distribution shift during training. This shift arises because the target values in Eq. (10) use actions sampled from the learned policy $\pi_{\theta}$, while the Q-function is trained only on actions sampled from the behavior policy that produced the dataset ($(\bm{s},\bm{a})\in\mathcal{D}$). Therefore, augmenting the standard Bellman operator with the advantage term $A(\bm{a^{\prime}}|\bm{s^{\prime}})$ can effectively mitigate estimation errors on OOD actions.
Proposition 3.3.
The advantage-based Bellman operator defined in Eq. (10) is $\gamma$-contractive with respect to the $L_{\infty}$ norm, and any initial $Q$-function can converge to a unique fixed point by repeatedly applying this operator.
Therefore, we define the unique fixed point of Eq. (10) as $Q_{\pi_{\theta}}^{A}$, and denote by $Q_{\pi_{\theta}}$ the true Q-function of $\pi_{\theta}$. The relationship between $Q_{\pi_{\theta}}^{A}$ and $Q_{\pi_{\theta}}$ is characterized in the following proposition.
Proposition 3.4.
The unique fixed point $Q_{\pi_{\theta}}^{A}$ of the advantage-based Bellman operator satisfies
$\|Q_{\pi_{\theta}}^{A}-Q_{\pi_{\theta}}\|_{\infty}\leq\frac{2\lambda R_{\rm max}}{(1-\gamma)^{2}}.$
Figure 2: Comparison of prior evaluation methods against our advantage-based evaluation. We visualize the Q-function and the policy for a fixed state. The thick blue solid line denotes the optimal value function. The solid line denotes the Q-function and the corresponding policy learned via advantage-based evaluation. The dashed line denotes the Q-function and policy learned via conservative evaluation. The dotted line denotes the Q-function and policy learned via standard Bellman backup evaluation.
Proposition 3.4 shows that the Q-function obtained by the advantage-based Bellman operator modifies the true Q-function derived from the standard Bellman operator within a bounded range of $2\lambda R_{\rm max}/(1-\gamma)^{2}$. We further illustrate the effectiveness of our advantage-based evaluation in Figure 2, where different line styles correspond to different evaluation methods. In the offline RL setting, due to the prohibition against interacting with the environment, directly applying standard Bellman backup (dotted line in Figure 2) results in an erroneous Q-function, which tends to overestimate the value of OOD actions and thus leads to an ineffective policy. Meanwhile, conservative evaluation (dashed line) indiscriminately penalizes all OOD actions, resulting in suppressed Q-values across these actions and limiting the learned policy to a sub-optimal policy near the support of the dataset. By contrast, our advantage-based evaluation (solid line) defines an advantage function that effectively modulates the Q-function obtained from Bellman backup, enabling the policy to discover optimal actions even beyond the support of the dataset. This phenomenon is further validated by our empirical results on the PointMaze tasks, as shown in Section 5.2.
4 Advantage-Based Diffusion Actor-Critic Algorithm
We now introduce our practical algorithm, Advantage-based Diffusion Actor-Critic (ADAC).
Diffusion Policy. We model our policy as the reverse process of a conditional diffusion model wang2022diffusion :
$\pi_{\theta}(\bm{a}\,|\,\bm{s})=p_{\theta}(\bm{a^{0:T}}\,|\,\bm{s})=\mathcal{N}(\bm{a^{T}};\bm{0},\bm{I})\prod_{i=1}^{T}p_{\theta}(\bm{a^{i-1}}\,|\,\bm{a^{i}},\bm{s}),$
where the terminal sample $a^{0}$ is used as the action for RL evaluation. During training, we sample $(\bm{s},\bm{a})$ pairs from the offline dataset $\mathcal{D}$ and construct noisy samples $\bm{a^{i}}=\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon}$ (Eq. (3)), where $\alpha_{i}=1-\beta_{i}$, $\bar{\alpha}_{i}=\prod_{j=1}^{i}\alpha_{j}$, and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. Following DDPM ho2020denoising , we train the following noise prediction model $\bm{\epsilon}_{\theta}(\bm{a^{i}},\bm{s},i)$ to approximate the added noise, which determines the reverse process $p_{\theta}(\bm{a^{i-1}}\,|\,\bm{a^{i}},\bm{s})$:
$\mathcal{L}_{d}(\theta)=\mathbb{E}_{i\sim\mathcal{U},\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}),(\bm{s},\bm{a})\sim\mathcal{D}}\left[\big{\|}\bm{\epsilon}-\bm{\epsilon}_{\theta}(\sqrt{\bar{\alpha}_{i}}\bm{a}+\sqrt{1-\bar{\alpha}_{i}}\bm{\epsilon},\bm{s},i)\big{\|}^{2}\right],$ (11)
where $\mathcal{U}$ denotes the uniform distribution over $\{1,\dots,T\}$. $\mathcal{L}_{d}(\theta)$ is a behavior cloning loss, and minimizing it enables the diffusion model to learn the behavior policy $\mu$. At inference time, an action $\bm{a^{0}}$ is generated by sampling $a^{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and iteratively applying the learned reverse process.
Algorithm 1 Advantage-Based Diffusion Actor-Critic
1:  Input Offline dataset $\mathcal{D}$, policy network $\pi_{\theta}$, critic networks $Q_{\phi}$
2:  Train a behavior policy $\mu$ by minimizing Eq. (11)
3:  Train a value function $V$ by minimizing Eq. (12)
4:  Train a transition model $P$ by minimizing Eq. (13)
5:  for each iteration do
6:     Obtain $A(\bm{a^{\prime}}|\bm{s^{\prime}})$ according to Eq. (9) // Advantage Calculation
7:     Update $\mathcal{L}_{q}(\phi)$ according to Eq. (14) // Critic Update
8:     Update $\mathcal{L}_{p}(\theta)$ according to Eq. (15)// Actor Update
9:     Soft update parameter $\phi$ and $\theta$// Target networks Update
10:  end for
Advantage Function. In our practical implementation, the value function is parameterized by a neural network with parameters $\varphi$ and trained by minimizing the following expectile regression loss:
$\mathcal{L}_{v}(\varphi)=\mathbb{E}_{(\bm{s},\bm{a},r,\bm{s^{\prime}})\sim\mathcal{D}}\big{[}L_{2}^{\tau}(r+\gamma V_{\varphi}(\bm{s^{\prime}})-V_{\varphi}(\bm{s}))\big{]}.$ (12)
We parameterize the transition dynamics using a neural network and learn a deterministic transition model, which we find sufficiently accurate and computationally efficient in practice. The model is trained by minimizing the following mean squared error (MSE) loss:
$\mathcal{L}_{p}(\psi)=\mathbb{E}_{(\bm{s},\bm{a},\bm{s^{\prime}})\sim\mathcal{D}}\big{[}\|P_{\psi}(\bm{s},\bm{a})-\bm{s^{\prime}}\|^{2}\big{]}.$ (13)
Therefore, the behavior policy $\mu$, the value function $V$, and the transition dynamics $P$ can be learned by minimizing Eq. (11), Eq. (12), and Eq. (13), respectively. The advantage function $A(\bm{a}\,|\,\bm{s})$ is then computed as defined in Eq. (9). All components are trained once using only offline data and remain fixed during the subsequent actor-critic updates, making this stage computationally inexpensive.
Actor-Critic. Following the advantage-based Bellman operator defined in Eq. (10), we define the loss for learning Q-function (critic) as:
$\mathcal{L}_{q}(\phi)=\mathbb{E}_{(\bm{s},\bm{a},\bm{s^{\prime}})\sim\mathcal{D},\bm{a^{\prime}}\sim\pi_{\theta}(\cdot\mid\bm{s^{\prime}})}\left[\left(r(\bm{s},\bm{a})+\gamma\big{(}Q_{\phi}(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}\mid\bm{s^{\prime}})\big{)}-Q_{\phi}(\bm{s},\bm{a})\right)^{2}\right].$ (14)
Here, the advantage function $A(\bm{a}\,|\,\bm{s})$ acts as an auxiliary correction term, learned once from offline data and held fixed during critic updates. To improve the policy, we incorporate Q-function guidance into the behavior cloning objective, encouraging the model to sample actions with greater estimated values. The resulting policy (actor) objective combines policy regularization and policy improvement:
$\mathcal{L}_{p}(\theta)=\mathcal{L}_{d}(\theta)-\alpha\mathbb{E}_{\bm{s}\sim\mathcal{D},\bm{a}\sim\pi_{\theta}}\left[Q_{\phi}(\bm{s},\bm{a})\right].$ (15)
We summarize our implementation in Algorithm 1. A key innovation of our method lies in the use of $A(\bm{a}\,|\,\bm{s})$; without this component, our approach degenerates to the DQL method wang2022diffusion , which uses standard Bellman backup. In addition to achieving new SOTA, we compare with DQL on PointMaze in Section 5.2 to visualize how advantage modulation facilitates superior OOD action selection.
5 Experiments
In this section, we begin by evaluating our method on the widely recognized D4RL benchmark fu2020d4rl . We then design a dedicated experiment on PointMaze—another task within D4RL—to empirically and visually demonstrate ADAC’s ability to identify beneficial OOD actions. Finally, we perform an ablation study to dissect the contribution of key components in our method.
Dataset. We evaluate our method on four distinct domains from the D4RL benchmark: Gym, AntMaze, Adroit, and Kitchen. The Gym-MuJoCo locomotion tasks are widely adopted and relatively straightforward due to their simplicity and dense reward signals. In contrast, AntMaze presents more challenging scenarios with sparse rewards, requiring the agent to compose suboptimal trajectories to reach long-horizon goals. The Adroit tasks, collected from human demonstrations, involve narrow state-action regions and demand strong regularization to ensure desired performance. Finally, the Kitchen environment poses a multi-task control problem where the agent must sequentially complete four sub-tasks, emphasizing long-term planning and generalization to unseen states.
Baseline. We consider a diverse array of baseline methods that exhibits strong results in each domain of tasks. For policy regularization-based method, we compare with the classic BC, TD3+BC fujimoto2021minimalist , BEARkumar2019stabilizing , BRACwu2019behavior , BCQfujimoto2019off , AWRpeng2019advantage , O-RLbrandfonbrener2021offline , and DQLwang2022diffusion . For pessimistic value function-based approach, we include CQLkumar2020conservative , IQLkostrikov2021offline , and REMagarwal2020optimistic . For model-based offline RL, we choose MoRelkidambi2020morel . For the classic online method, we include SAChaarnoja2018soft . For conditional sequence modeling approaches, we include DTchen2021decision , Diffuserjanner2022planning , and DDajay2022conditional . We report the performance of baseline methods either from the best results published in their respective papers or from wang2022diffusion .
Table 1: Normalized average returns on D4RL tasks, averaged over the final 10 evaluations across 4 seeds.
Gym Tasks BC TD3+BC CQL IQL MoRel DT Diffuser DD DQL ADAC (Ours) halfcheetah-medium 42.6 48.3 44.0 47.4 42.1 42.6 44.2 49.1 51.1 58.0$\pm$0.3 hopper-medium 52.9 59.3 58.5 66.3 95.4 67.6 58.5 79.3 90.5 93.5$\pm$4.2 walker2d-medium 75.3 83.7 72.5 78.3 77.8 74.0 79.7 82.5 87.0 87.6$\pm$2.0 halfcheetah-medium-replay 36.6 44.6 45.5 44.2 40.2 36.6 42.2 39.3 47.8 52.5$\pm$0.8 hopper-medium-replay 18.1 60.9 95.0 94.7 93.6 82.7 96.8 100.0 101.3 102.1$\pm$1.1 walker2d-medium-replay 26.0 81.8 77.2 73.9 49.8 66.6 61.2 75.0 95.5 96.0$\pm$1.6 halfcheetah-medium-expert 55.2 90.7 91.6 86.7 53.3 86.8 79.8 90.6 96.8 106.1$\pm$1.0 hopper-medium-expert 52.5 98.0 105.4 91.5 108.7 107.6 107.2 111.8 111.1 112.5$\pm$1.0 walker2d-medium-expert 107.5 110.1 108.8 109.6 95.6 108.1 108.4 108.8 110.1 112.3$\pm$0.9 Average 51.9 75.3 77.6 77.0 72.9 74.7 75.3 81.8 88.0 91.2 AntMaze Tasks BC TD3+BC CQL IQL BEAR DT BCQ O-RL DQL ADAC (Ours) antmaze-umaze 54.6 78.6 74.0 87.5 73.0 59.2 78.9 64.3 93.4 98.2$\pm$4.5 antmaze-umaze-diverse 45.6 71.4 84.0 62.2 61.0 53.0 55.0 60.7 66.2 76.0$\pm$9.9 antmaze-medium-play 0.0 10.6 61.2 71.2 0.0 0.0 0.0 0.3 76.6 86.5$\pm$9.8 antmaze-medium-diverse 0.0 3.0 53.7 70.0 8.0 0.0 0.0 0.0 78.6 88.7$\pm$10.2 antmaze-large-play 0.0 0.2 15.8 39.6 6.7 0.0 6.7 0.0 46.4 69.8$\pm$12.4 antmaze-large-diverse 0.0 0.0 14.9 47.5 2.2 0.0 2.2 0.0 56.6 64.6$\pm$12.7 Average 16.7 27.3 50.6 63.0 23.7 18.7 23.8 20.9 69.6 80.6 Adroit Tasks BC BRAC-v CQL IQL BEAR REM BCQ SAC DQL ADAC (Ours) pen-human 25.8 0.6 35.2 71.5 -1.0 5.4 68.9 4.3 72.8 74.4$\pm$18.6 pen-cloned 38.3 -2.5 27.2 37.3 26.5 -1.0 44.0 -0.8 57.3 80.5$\pm$14.3 Average 32.1 -1.0 31.2 54.4 12.8 2.2 56.5 1.8 65.1 77.5 Kitchen Tasks BC BRAC-v CQL IQL BEAR AWR BCQ SAC DQL ADAC (Ours) kitchen-complete 33.8 0.0 43.8 62.5 0.0 0.0 8.1 15.0 84.0 87.9$\pm$6.7 kitchen-partial 33.8 0.0 49.8 46.3 13.1 15.4 18.9 0.0 60.5 65.2$\pm$7.0 kitchen-mixed 47.5 0.0 51.0 51.0 47.2 10.6 8.1 2.5 62.6 68.3$\pm$5.8 Average 38.4 0.0 48.2 53.3 20.1 8.7 11.7 5.8 69.0 73.8
5.1 Benchmark Results
Our method is evaluated on four task domains, with results summarized in Table 1. We also provide domain-specific analysis to highlight key performance characteristics.
Results for Gym Tasks. While prior methods already achieve strong performance on Gym tasks, ADAC consistently delivers further improvements. Among all Gym tasks, halfcheetah is notably the most challenging, often yielding the lowest scores. However, ADAC achieves the largest performance boost on this task family, with an improvement of 10%. According to the protocol established in fu2020d4rl , a normalized score of 100 corresponds to an expert-level policy. Therefore, achieving an average normalized score exceeding 90 demonstrates the practical effectiveness of our method.
Figure 3: Performance comparison of DQL and ADAC on three AntMaze tasks: Umaze, Medium Diverse, and Large Play. Each method was trained with 4 random seeds, and the reward curves were smoothed with a running average ($n=10$). In this figure, the solid lines correspond to the mean and the shaded regions correspond to the standard deviation.
Results for AntMaze Tasks. The AntMaze tasks are particularly challenging due to the sparse rewards and the prevalence of sub-optimal trajectories. Nevertheless, this is precisely where ADAC proves advantageous, as exploring OOD actions becomes essential for achieving high performance under such circumstances. Remarkably, we achieve over 15% improvements across all baseline methods. We visualize the learning process of the pre-SOTA method DQL and ADAC in Figure 3. As illustrated in Figure 3, ADAC consistently outperforms DQL in terms of both performance and stability across all three AntMaze task variants.
Results for Adroit and Kitchen Tasks. In the Adroit tasks, the narrow support of human demonstrations leads to substantial extrapolation error. While both DQL and ADAC incorporate policy regularization to address this issue, our advantage-based approach further mitigates it by effectively handling unseen actions beyond the demonstrations. As a result, we achieve a 20% performance improvement over existing baselines. Additionally, the Kitchen tasks require long-horizon value optimization, and we observe that ADAC also outperforms in this domain.
5.2 Visualizing OOD Action Selection in PointMaze
Figure 4: Sparse reward PointMaze: dataset and method performance. Top: 853 sub-optimal trajectories are gathered with only terminal reward. Three trajectory patterns—Left (33%), Middle (22%), Right (45%)—span lengths of 200 to 1000 steps (the optimal length is 142). Bottom: The first two subfigures illustrate the trajectories generated by DQL and ADAC after training on the dataset, respectively. The last subfigure summarizes the distribution of trajectory lengths for the offline dataset, DQL, and ADAC.
In the previous subsection, we demonstrated that our method achieves state-of-the-art performance across a wide range of D4RL benchmark tasks, with particularly large gains on challenging sparse-reward environments. This improvement can be largely attributed to our newly designed advantage function, which enables the selection of beneficial OOD actions—a capability especially critical in sparse-reward tasks, where exploiting such actions is essential for high returns.
To provide further visual evidence that the advantage function guides the selection of beneficial OOD actions, we conduct a comparative experiment on PointMaze. Specifically, we construct a toy environment based on the latest gymnasium-robotics implementation of PointMaze, derived from the Maze2D environment in the D4RL suite. As shown in Figure 4, the green circle indicates the agent’s starting position, the red star denotes the goal, and the beige squares represent static obstacles. The task involves navigating a 2-DoF point agent through a maze with obstacles to a fixed goal using Cartesian (x, y) actuation. This is a sparse-reward task: the agent receives a reward of 1 only upon reaching the goal and zero at all other steps.
We manually collected 853 successful yet sub-optimal trajectories to construct the offline training dataset, each shorter than 1000 steps but significantly longer than the shortest length of approximately 142 steps, which corresponds to a straight-line path from the start to the goal. As shown in the last subfigure of Figure 4, the offline dataset includes no optimal trajectories (length $<$ 175), only 3.4% near-optimal (length $<$ 225), and 19.1% competitive trajectories (length $<$ 300). These trajectories fall into three distinct modes—Left, Middle, and Right routes—depending on which corridor the agent takes to bypass the obstacles (see corresponding subfigures in Figure 4). The dataset comprises 391 391 Q-learning-style tuples of the form $(\bm{s},\bm{a},\bm{s^{\prime}},r,\text{done})$. The collection strategy and detailed trajectory quality analysis are provided in Appendix C.8.
We train both DQL wang2022diffusion and ADAC for 50 000 steps using the constructed dataset to obtain their respective policies. Each policy is then evaluated in our environment, generating 300 trajectories per method, as illustrated in the bottom row of Figure 4.
Compared to the full set of offline trajectories, the ADAC-generated trajectories contain a substantial number of straight-line paths from the start to the goal—routes that are entirely absent in the dataset. This visual evidence strongly supports that ADAC is capable of identifying and selecting superior OOD actions. In contrast, the trajectories generated by DQL, while showing moderate improvement over the offline data, still heavily reflect left, middle, and right patterns, indicating strong behavior cloning tendencies. Since the key distinction between DQL and ADAC lies in the introduction of advantage modulation, these visualizations clearly validate the effectiveness of the advantage function in enabling the selection of beneficial OOD actions. The last subfigure of Figure 4 provides quantitative evidence that ADAC substantially outperforms DQL in trajectory quality.
5.3 Ablation Study
Figure 5: Ablation studies on the effect of $\kappa$, the quantile defining the threshold for positive advantage, in AntMaze and Gym Locomotion domains. Higher values of $\kappa$ (approaching 1) correspond to using the value of the optimal action under the behavior policy as the threshold, while lower $\kappa$ values relax this criterion. All results are averaged over 4 independent random seeds.
The key innovation of our method lies in the design of the advantage function defined in Eq. (9). This formulation introduces a parameter $\kappa$ to control the level of conservatism in our algorithm. We investigate how $\kappa$ interacts with key factors including dataset quality, task difficulty, and reward sparsity. This investigation aims to provide guidance for tuning ADAC when applying it to new tasks. We conduct ablation studies on six sparse-reward AntMaze environments with three levels of difficulty. We also evaluate on nine Gym Locomotion tasks under three dataset quality settings. As shown in Fig. 5, we vary $\kappa$ over the range $\{0.55,0.65,0.75,0.85,0.95\}$ to assess its impact.
We find that in sparse-reward AntMaze tasks, smaller values of $\kappa$ (e.g., 0.55 and 0.65) yield superior performance, whereas in dense-reward Gym tasks, larger values (e.g., 0.75) lead to better results. This observation aligns with the intended design of our method: smaller $\kappa$ values promote the selection of OOD actions, which is essential for achieving high returns in sparse-reward settings.
One potential limitation of our method is the need to sample multiple actions ($N\equiv 25$ in this paper) from the behavior policy to compute the advantage function. However, our ablation study shows that in each task domain, competitive performance can be achieved across at least three different values of $\kappa$, indicating that the algorithm is relatively insensitive to $\kappa$. This suggests that the number of candidate actions can be reduced without significantly affecting performance. We implement our algorithm using the JAX/Flax framework, which offers faster training and inference speed than the DQL method and is comparable to the optimized EDP kang2023efficient implementation as shown in Figure 9.
6 Conclusion
In this work, we propose ADAC, a novel offline RL algorithm that systematically evaluates the quality of OOD actions to balance conservatism and generalization. ADAC represents a pioneering attempt to explicitly assess OOD actions and to selectively encourage beneficial ones, while discouraging risky ones to maintain conservatism. We validate the effectiveness of advantage modulation through a series of custom PointMaze experiments and demonstrate state-of-the-art performance across almost all tasks in the D4RL benchmark. The empirical results further indicate that ADAC is particularly effective in more challenging tasks.
Supplementary Material
appendix.Aappendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4subsection.C.5subsection.C.6subsection.C.7subsection.C.8
Appendix A Related Work
Offline RL focuses on learning effective policies solely from a pre-collected behavior dataset and has demonstrated significant success in practical applications [29, 31, 24]. The existing literature on offline RL can be classified into four main categories:
Pessimistic value-based methods achieve conservatism by incorporating penalty terms into the value optimization objective, discouraging the value function from being overly optimistic on out-of-distribution (OOD) actions. Specifically, CQL [21] applies equal penalization to Q-values for all OOD samples, whereas EDAC [3] and PBRL [4] adjust the penalization based on the uncertainty level of the Q-value, measured using a neural network ensemble.
Regularized policy-based methods constrain the learned policy to stay close to the behavior policy, thereby avoiding OOD actions. For instance, BEAR [20] constrains the optimized policy by minimizing the MMD distance to the behavior policy. BCQ [11] restricts the action space to those present in the dataset by utilizing a learned Conditional-VAE (CVAE) behavior-cloning model. Alternatively, TD3+BC [10] simply adds a behavioral cloning regularization term to the policy optimization objective and achieves excellent performance across various tasks. IQL [19] adopts an advantage-weighted behavior cloning approach, learning Q-value functions directly from the dataset. Meanwhile, DQL [36] leverages diffusion policies as an expressive policy class to enhance behavior-cloning. Our work falls into this category as we also incorporate a behavior-cloning term.
Conditional sequence modeling methods induce conservatism by limiting the policy to replicate behaviors from the offline dataset [7, 38]. This approach naturally follows a supervised learning paradigm. Additionally, trajectories can be formulated as conditioned generative models and generated by diffusion models that satisfy conditioned constraints [14, 2].
Model-based methods incorporate conservatism to prevent the policy from overgeneralizing to regions where the dynamics model predictions are unreliable. For example, COMBO [39] extends CQL to a model-based setting by enforcing small Q-values for OOD samples generated by the dynamics model. RAMBO [30] incorporates conservatism by adversarially training the dynamics model to minimize the value function while maintaining accurate transition predictions. Most model-based methods achieve conservatism through uncertainty quantification, penalizing rewards in regions with high uncertainty. Specifically, MOPO [40] uses the max-aleatoric uncertainty quantifier, MOReL [18] employs the max-pairwise-diff uncertainty quantifier, and MOBILE [34] leverages the Model-Bellman inconsistency uncertainty quantifier. Recently, [8] achieves conservatism by incorporating the value function inconsistency loss, enabling the training of a more reliable model.
Appendix B Proof of Propositions
In this subsection, we provide comprehensive and complete proofs for our propositions listed in Section 3.
Proof of Proposition 3.1.
Proof.
Denote $\delta(\bm{s},\bm{a},\bm{s^{\prime}})=r(\bm{s},\bm{a})+\gamma V(\bm{s^{\prime}})-V(\bm{s})$, the regression problem in Eq. (5) can be rewritten as
$\mathcal{L}(V)=\mathbb{E}_{(\bm{s},\bm{a},r,\bm{s^{\prime}})\sim\mathcal{D}}\big{[}L_{2}^{\tau}(\delta(\bm{s},\bm{a},\bm{s^{\prime}}))\big{]}.$
To find the optimal $V(\bm{s})$, we take the derivative of $\mathcal{L}(V)$ with respect to $V(\bm{s})$ conditioning on $\bm{s}$:
$\frac{\partial\mathcal{L}(V)}{\partial V(\bm{s})}=$ $\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bigg{[}\frac{\partial L_{2}^{\tau}(\delta)}{\partial V(\bm{s})}\bigg{]}$
=\ = $\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bigg{[}\frac{\partial L_{2}^{\tau}(\delta)}{\partial\delta}\cdot\frac{\partial\delta}{\partial V(\bm{s})}\bigg{]}$
=\ = $\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}2|1-\mathds{1}(\delta<0)|\delta(\bm{s},\bm{a},\bm{s^{\prime}})\cdot(-1)\big{]},$
where the exchange of partial derivative and expectation is due to dominated convergence theorem since both $r$ and $V$ are bounded.
From the fact that the solution $V_{\tau}(\bm{s})$ satisfies $\frac{\partial\mathcal{L}(V)}{\partial V(\bm{s})}\big{|}_{V_{\tau}(\bm{s})}=0$, we get
$\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}|\tau-\mathds{1}(r(\bm{s},\bm{a})+\gamma V_{\tau}(\bm{s^{\prime}})-V_{\tau}(\bm{s})<0)|(r(\bm{s},\bm{a})+\gamma V_{\tau}(\bm{s^{\prime}})-V_{\tau}(\bm{s}))\big{]}=0.$
In expectile regression, the $\tau$-expectile $\mu_{\tau}$ of a random variable $X$ satisfies
$\mathbb{E}\big{[}|\tau-\mathds{1}(X-\mu_{\tau}<0)|(Y-\mu_{\tau})\big{]}=0.$
As a result, this implies that the solution $V_{\tau}(\bm{s})$ is the $\tau$-expectile of the target $r(\bm{s},\bm{a})+\gamma V_{\tau}(\bm{s^{\prime}})$. Therefore, we conclude that
$V_{\tau}(\bm{s})=\mathbb{E}^{\tau}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}r(\bm{s},\bm{a})+\gamma V_{\tau}(\bm{s^{\prime}})\big{]},$
which finish our proof. ∎
Proof of Proposition 3.2.
Proof.
Define the $\tau$-expectile Bellman operator as
$\mathcal{T_{\tau}}V(\bm{s}):=\mathbb{E}^{\tau}_{\bm{a}\sim\mu(\cdot\,|\,\bm{s}),\bm{s^{\prime}}\sim P(\cdot\,|\,\bm{s},\bm{a})}\big{[}r(\bm{s},\bm{a})+\gamma V(\bm{s^{\prime}})\big{]}.$
From Eq. (6), we know that $\mathcal{T}_{\tau}V_{\tau}(\bm{s})=V_{\tau}(\bm{s})$, which means $V_{\tau}(\bm{s})$ is a fixed point for $\tau$-expectile Bellman operator $\mathcal{T}_{\tau}$.
Suppose there is another fixed point $W_{\tau}(\bm{s})$ for $\mathcal{T}_{\tau}$. It holds that
$\|V_{\tau}-W_{\tau}\|_{\infty}=\|\mathcal{T}_{\tau}V_{\tau}-\mathcal{T}_{\tau}W_{\tau}\|_{\infty}\leq\gamma\|V_{\tau}-W_{\tau}\|_{\infty},$
which means that $V_{\tau}=W_{\tau}$. Therefore, we have shown that $V_{\tau}$ is the unique fixed point for $\mathcal{T}_{\tau}$.
For $\tau_{1}\leq\tau_{2}$ and a bounded random variable $X$, we have
$\mathbb{E}^{\tau_{1}}[X]\leq\mathbb{E}^{\tau_{2}}[X].$
As a result, we get $\mathcal{T}_{\tau_{1}}V\leq\mathcal{T}_{\tau_{2}}V$. Therefore, for its fixed point, we have $V_{\tau_{1}}\leq V_{\tau_{2}}$.
It can be shown that
$V_{\tau}(\bm{s})=\mathcal{T}_{\tau}V_{\tau}(\bm{s})\leq\max\limits_{\bm{a}\in\rm\mu(\cdot|\bm{s})}\big{[}r(\bm{s},\bm{a})+\gamma\|V_{\tau}\|_{\infty}\big{]}\leq\frac{2R_{\rm max}}{1-\gamma}.$
Therefore, we have demonstrated that $V_{\tau}(\bm{s})$ is bounded and monotonically non-decreaing in $\tau$. Consequently, there exists a limit $\bar{V}(\bm{s})$ such that
$\lim\limits_{\tau\to 1}V_{\tau}(\bm{s})=\bar{V}(\bm{s}).$
Define a random variable
$X^{\tau}=r(\bm{s},\bm{A})+\gamma V_{\tau}(\bm{S^{\prime}}),\quad\bm{A}\sim\mu(\cdot\,|\,\bm{s}),\bm{S^{\prime}}\sim P(\cdot\,|\,\bm{s},\bm{A}),$
and its limit $\bar{X}=r(\bm{s},\bm{A})+\gamma\bar{V}(\bm{S^{\prime}})$. It follows that
$\lim\limits_{\tau\to 1}V_{\tau}(\bm{s})=\lim\limits_{\tau\to 1}\mathbb{E}^{\tau}[X^{\tau}]\overset{(1)}{=}\lim\limits_{\tau\to 1}\mathbb{E}^{\tau}[\bar{X}],$
where (1) comes from
$|\mathbb{E}^{\tau}[X^{\tau}]-\mathbb{E}[\bar{X}]|\leq\mathbb{E}|X^{\tau}-\bar{X}|.$
From Lemma 1 in [19], which states that
$\lim\limits_{\tau\to 1}\mathbb{E}^{\tau}[\bar{X}]=\max(\bar{X}).$
Therefore, we get
$\bar{V}(\bm{s})=\max\limits_{\bm{a}\in\mu(\cdot|\bm{s})}\big{[}r(\bm{s},\bm{a})+\gamma\max\limits_{\bm{s^{\prime}}\sim P}\bar{V}(\bm{s})\big{]}.$
For deterministic transition probability $P$, we have
$\bar{V}(\bm{s})=\max\limits_{\bm{a}\in\mu(\cdot|\bm{s})}\big{[}r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bar{V}(\bm{s})\big{]}.$
Define the batch-optimal Bellman operator as
$\mathcal{T}^{\ast}_{\mu}V(\bm{s})=\max\limits_{\bm{a}\in\mu(\cdot|\bm{s})}\big{[}r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}V(\bm{s})\big{]}.$
It follows that $\bar{V}(\bm{s})$ and $V^{\ast}_{\mu}(\bm{s})$ are both fixed point for $\mathcal{T}^{\ast}_{\mu}$. By a similar argument for $\mathcal{T}_{\tau}$, we know that $\mathcal{T}^{\ast}_{\mu}$ is $\gamma$-contractive and has a unique fixed point. As a result, it holds that
$\lim\limits_{\tau\to 1}V_{\tau}(\bm{s})=V_{\mu}^{\ast}(\bm{s})$
for a deterministic transition probability. Overall, we finish our proof. ∎
Proof of Proposition 3.3.
Proof.
Let $Q_{1}$ and $Q_{2}$ be two arbitrary $Q$-functions. We have
$\|\mathcal{T}_{A}^{\pi_{\theta}}Q_{1}-\mathcal{T}_{A}^{\pi_{\theta}}Q_{2}\|_{\infty}=$ $\max\limits_{\bm{s},\bm{a}}\big{|}r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\big{[}Q_{1}(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}|\bm{s^{\prime}})\big{]}$
$-r(\bm{s},\bm{a})-\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\big{[}Q_{2}(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}|\bm{s^{\prime}})\big{]}\big{|}$
$=$ $\max\limits_{\bm{s},\bm{a}}\big{|}\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\big{[}Q_{1}(\bm{s^{\prime}},\bm{a^{\prime}})-Q_{2}(\bm{s^{\prime}},\bm{a^{\prime}})\big{]}\big{|}$
$\leq$ $\gamma\max_{\bm{s},\bm{a}}\|Q_{1}-Q_{2}\|_{\infty}$
$=$ $\gamma\|Q_{1}-Q_{2}\|_{\infty}.$
Therefore, $\mathcal{T}_{A}^{\pi_{\theta}}$ is a $\gamma$-contraction operator which naturally implies any initial Q-function can converge to a unique fixed point by repeatedly applying this operator. ∎
Proof of Proposition 3.4.
Proof.
We first show that $\|A(\bm{a^{\prime}}|\bm{s^{\prime}})\|_{\infty}\leq 2R_{\rm max}/(1-\gamma)$. From the proof of Proposition 3.2, we know that
$\mathcal{T_{\tau}}V(\bm{s}):=\mathbb{E}^{\tau}_{\bm{a}\sim\mu(\cdot\,|\,\bm{s}),\bm{s^{\prime}}\sim P(\cdot\,|\,\bm{s},\bm{a})}\big{[}r(\bm{s},\bm{a})+\gamma V(\bm{s^{\prime}})\big{]}.$
Since the $\tau$-expectile of a random variable cannot exceed its maximum, for any $V$, we have
$\|\mathcal{T}_{\tau}V\|_{\infty}\leq R_{\rm max}+\gamma\|V\|_{\infty}.$
From Eq. (6), we know that
$\|V_{\tau}\|_{\infty}=\|\mathcal{T}_{\tau}V_{\tau}\|_{\infty}\leq R_{\rm max}+\gamma\|V_{\tau}\|_{\infty}.$
It follows that
$\|V_{\tau}\|_{\infty}\leq\frac{R_{\rm max}}{1-\gamma},\forall\tau.$
Therefore, we get $\|V\|_{\infty}\leq R_{\rm max}/(1-\gamma)$ and $\|A(\bm{a^{\prime}}|\bm{s^{\prime}})\|_{\infty}\leq 2R_{\rm max}/(1-\gamma)$.
From the advantage-based operator $\mathcal{T}^{\pi_{\theta}}_{A}$, we have
$\mathcal{T}^{\pi_{\theta}}_{A}Q(\bm{s},\bm{a})=$ $\ r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\bigg{[}Q(\bm{s^{\prime}},\bm{a^{\prime}})+\lambda A(\bm{a^{\prime}}|\bm{s^{\prime}})\bigg{]}$
$=$ $\ r(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\bigg{[}Q(\bm{s^{\prime}},\bm{a^{\prime}})\bigg{]}+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\big{[}\lambda A(\bm{a^{\prime}}|\bm{s^{\prime}})\big{]}$
$=$ $\ \mathcal{T}^{\pi_{\theta}}Q(\bm{s},\bm{a})+\gamma\mathbb{E}_{\bm{s^{\prime}}\sim P,\bm{a^{\prime}}\sim\pi_{\theta}}\big{[}\lambda A(\bm{a^{\prime}}|\bm{s^{\prime}})\big{]},$
where $\mathcal{T}^{\pi_{\theta}}$ is the standard Bellman operator. From the boundedness of $A(\bm{a}|\bm{s})$, we have
$\mathcal{T}^{\pi_{\theta}}Q(\bm{s},\bm{a})-\gamma\frac{2\lambda R_{\rm max}}{1-\gamma}\leq\mathcal{T}^{\pi_{\theta}}_{A}Q(\bm{s},\bm{a})\leq\mathcal{T}^{\pi_{\theta}}Q(\bm{s},\bm{a})+\gamma\frac{2\lambda R_{\rm max}}{1-\gamma}.$
Iteratively applying this operator to obtain the fixed point, we get
$Q_{\pi_{\theta}}-\frac{2\lambda R_{\rm max}}{(1-\gamma)^{2}}\leq Q_{\pi_{\theta}}^{A}\leq Q_{\pi_{\theta}}+\frac{2\lambda R_{\rm max}}{(1-\gamma)^{2}},\forall\bm{s},\bm{a},$
which implies our conclusion. ∎
Appendix C Experimental Details
C.1 Advantage Function Characterization and Regularization
To provide an overview of the learned advantage function, we report summary statistics computed across 20 D4RL tasks using the best-performing hyperparameter setting (Appendix C.4). For each task, we aggregate advantage values from four independent training runs and report: (1) the mean and standard deviation of both positive and negative advantages, and (2) the proportion of samples exhibiting positive advantages.
The results in Table 2 reveal notable variability in the distribution of advantage values across tasks. In particular, the proportion of positive advantages differs substantially between environments, reflecting how often the learned value function favors alternative actions over those observed in the dataset. We observe that some tasks display a low proportion of positive advantages—reflecting fewer opportunities for improvement—whereas others show substantially higher positive ratios, indicating greater diversity in action quality and more room for enhancement. These statistics are determined by factors such as the quality of the behavior policy, hyperparameters like $\kappa$, and the learned value function $V$, among others.
Table 2: Advantage statistics for 20 D4RL tasks.
Task Name Positive Negative Pos. (%)
halfcheetah-medium $1.62_{\pm 0.2}$ $-0.11_{\pm 0.3}$ 32.8
halfcheetah-medium-replay $1.84_{\pm 0.3}$ $-0.90_{\pm 0.0}$ 30.4
halfcheetah-medium-expert $2.00_{\pm 0.2}$ $-1.75_{\pm 0.3}$ 38.7
hopper-medium $0.48_{\pm 0.1}$ $-0.14_{\pm 0.0}$ 22.1
hopper-medium-replay $1.25_{\pm 0.2}$ $-0.83_{\pm 0.1}$ 38.3
hopper-medium-expert $0.39_{\pm 0.2}$ $-0.55_{\pm 0.0}$ 10.1
walker2d-medium $0.70_{\pm 0.1}$ $-0.06_{\pm 0.0}$ 43.5
walker2d-medium-replay $1.87_{\pm 0.3}$ $-0.13_{\pm 0.0}$ 20.0
walker2d-medium-expert $2.33_{\pm 0.3}$ $-2.20_{\pm 0.5}$ 26.2
antmaze-umaze $1.70_{\pm 0.2}$ $-1.05_{\pm 0.1}$ 52.8
antmaze-umaze-diverse $0.03_{\pm 0.0}$ $-0.04_{\pm 0.0}$ 34.0
antmaze-medium-play $0.58_{\pm 0.1}$ $-0.40_{\pm 0.1}$ 44.7
antmaze-medium-diverse $0.48_{\pm 0.05}$ $-0.26_{\pm 0.05}$ 54.0
antmaze-large-play $0.58_{\pm 0.02}$ $-0.28_{\pm 0.05}$ 48.3
antmaze-large-diverse $0.42_{\pm 0.08}$ $-0.23_{\pm 0.02}$ 62.3
pen-human $2.34_{\pm 1.1}$ $-1.74_{\pm 0.6}$ 41.3
pen-cloned $1.14_{\pm 0.2}$ $-1.01_{\pm 0.1}$ 33.4
kitchen-complete $1.15_{\pm 0.2}$ $-0.85_{\pm 0.1}$ 42.3
kitchen-partial $0.47_{\pm 0.1}$ $-0.44_{\pm 0.2}$ 31.1
kitchen-mixed $0.52_{\pm 0.0}$ $-0.74_{\pm 0.0}$ 33.8
Advantage Soft Clipping.
Figure 6: Visualization of the $\text{softclip}(x)$.
To prevent unstable learning dynamics caused by extreme advantage values, we apply a soft clipping transformation to all computed advantages. The function is defined as
$\text{softclip}(x)=\begin{cases}\lambda_{p}\cdot{\tanh\bigl{(}\frac{x}{\lambda_{p}}\bigr{)}},&x\geq 0,\\[6.0pt] \lambda_{n}\cdot{\tanh\bigl{(}\frac{x}{\lambda_{n}}\bigr{)}},&x<0,\end{cases}$
where $\lambda_{p}$ and $\lambda_{n}$ serve as scaling factors for positive and negative values, respectively, replacing the single $\lambda$ used in Eq. (10) to enhance empirical performance in our implementation.
This formulation softly bounds the advantage values, with smooth saturation in the tails and a near-linear response around zero. Unlike hard clipping, it avoids sharp discontinuities while preserving the relative differences between actions, which is important for effective Q-function learning. A visualization of the softclip transformation is shown in Figure 6.
In practice, we observe that performance is largely insensitive to the precise values of $\lambda_{p}$ and $\lambda_{n}$, as long as their ratio is maintained. Specifically, setting $\lambda_{p}$ to approximately $1.5\times\lambda_{n}$ (e.g., such as $\lambda_{p}=6$ and $\lambda_{n}=4$) consistently yields stable gradients and expressive advantage signals. These results indicate that the method is robust to the specific choice of these hyperparameters, provided the relative scaling is preserved.
C.2 Diffusion Actor and Residual-Critic Architecture Design
Our method jointly optimizes three network modules during main training: a diffusion-based actor, a Q-function critic, and a value function $V$. This section describes the architecture of these components, excluding auxiliary networks used for advantage computation.
Diffusion Actor.
The actor is instantiated as a denoising diffusion probabilistic model (DDPM) with a variance-preserving (VP) noise schedule. The noise predictor is a five-layer multilayer perceptron (MLP) with Mish activations. We set the number of denoising steps to 10 across all tasks, balancing expressiveness with computational efficiency.
Critic Networks.
Both the Q-function and value function are instantiated in two variants: a standard MLP and a residual architecture comprising 16 residual blocks. We observe that the residual networks significantly improve both training stability and final performance in complex tasks, particularly in sparse-reward environments such as AntMaze. We attribute this to the increased representational capacity of the residual architecture, which is better suited to capturing the fine-grained structure of value functions when learning targets are noisy or heterogeneous. In contrast, shallow MLPs tend to underfit in such regimes, leading to unstable or overly conservative estimates. An illustration of the residual block is provided in Figure 7.
Figure 7: Residual block architecture used in critic networks.
C.3 Repeated Sampling for Value-Guided Diffusion Policies
Diffusion-based policies are capable of modeling expressive, multimodal action distributions conditioned on state. However, this flexibility introduces sampling stochasticity, where single-sample rollouts may fall into suboptimal modes. To address this, we adopt a repeated sampling strategy guided by the learned Q-function, which enhances both training stability and evaluation performance by enabling more informed action selection.
Training-Time Sampling: Max-Q Backup.
During critic updates, we widely employ a Max-Q Backup mechanism from CQL [21]: for each transition, multiple candidate actions are sampled from the policy at the next state, and the Q-target is computed using the maximum or softmax-weighted Q-value among them. This mitigates underestimation bias caused by poor single-sample rollouts and reduces the variance of the TD targets. We observe that modest sample counts (e.g., 3–5) already improve stability, while more complex tasks—such as halfcheetah-medium-replay and antmaze-medium-diverse benefit from larger sample sizes (e.g., 10). As shown in Figure 8, increasing the number of backup samples leads to higher predicted Q-values and improved empirical returns.
(a) Episode returns under varying backup counts.
(b) Training-Time average Q value per batch.
Figure 8: Impact of repeated sampling in both training and inference stages in halfcheetah-medium-replay.
Evaluation-Time Sampling: Q-Guided Inference.
At evaluation time, we sample a large set of candidate actions (typically 50–200) from the diffusion model, and select actions via Q-weighted sampling. Specifically, Q-values are transformed into a softmax distribution, from which the final action is drawn. This Q-guided inference biases the policy toward high-value modes while retaining stochasticity, offering a principled alternative to greedy or deterministic selection. Across tasks, we consistently observe superior returns compared to single-sample decoding.
This effect stems from the multimodal nature of diffusion policies: only a subset of modes yield high return, especially in sparse-reward settings. Without repeated sampling, the critic may overlook these high-reward regions, leading to pessimistic target estimates and suboptimal updates. Expanding the candidate set increases the likelihood of capturing valuable modes, thereby improving both value estimation and policy quality.
C.4 Hyperparameter Setup
We highlight several key components of our approach, including the use of the $\kappa$ quantile to define reward thresholds, the integration of residual blocks to enhance the expressiveness of the critic, and repeated sampling to exploit the multimodal capacity of the diffusion model. These components play a central role in achieving strong performance across tasks. Table 3 provides the full set of hyperparameters; all other parameters follow default configurations from DQL [36] without further modification.
Table 3: Hyperparameter configurations for all evaluated tasks. We report the quantile threshold $\kappa$, a boolean indicating whether max Q backups are applied, the number of backup times $n$, and the backbone used for the critic network.
Task $\boldsymbol{\kappa}$ Max Q Backup $\boldsymbol{n}$ Critic Net
halfcheetah-medium 0.75 True 5 ResNet
halfcheetah-medium-replay 0.75 True 5 ResNet
halfcheetah-medium-expert 0.75 True 10 ResNet
hopper-medium 0.75 False 1 ResNet
hopper-medium-replay 0.75 True 5 ResNet
hopper-medium-expert 0.95 False 1 ResNet
walker2d-medium 0.65 True 3 ResNet
walker2d-medium-replay 0.85 False 1 ResNet
walker2d-medium-expert 0.75 False 1 MLP
antmaze-umaze 0.55 True 10 ResNet
antmaze-umaze-diverse 0.65 True 10 ResNet
antmaze-medium-play 0.65 True 10 ResNet
antmaze-medium-diverse 0.65 True 10 ResNet
antmaze-large-play 0.65 True 10 ResNet
antmaze-large-diverse 0.55 True 10 ResNet
pen-human 0.65 True 3 MLP
pen-cloned 0.65 True 3 MLP
kitchen-complete 0.65 False 1 MLP
kitchen-partial 0.65 False 1 MLP
kitchen-mixed 0.65 False 1 MLP
C.5 Alternative Implementation
While expectile regression provides a theoretically grounded approach for estimating the optimal value function from offline data, we identify an IQL-style alternative that achieves comparable empirical performance and is easier to tune, particularly under highly suboptimal dataset quality and coverage. This approach couples value estimation with policy optimization, leveraging the evolving policy to gradually shift the data distribution toward regions with more informative value estimates. As a result, it often makes hyperparameter tuning, such as selecting $\tau$, more robust in practice.
Algorithm 2 Alternative Impl. for ADAC
1:  Input initial all parameters
2:  Train a diffusion behavior policy
3:  Train a MLP transition model $P$
4:  for each iteration do
5:     Update $\mathcal{L}_{v}(\varphi)$ according to Eq. (16)
6:     Obtain $A(\bm{a^{\prime}}|\bm{s^{\prime}})$ according to Eq. (9)
7:     Update $\mathcal{L}_{q}(\phi)$ according to Eq. (14)
8:     Update $\mathcal{L}_{p}(\theta)$ according to Eq. (15)
9:     Soft update parameter $\phi$ and $\theta$
10:  end for
Building on this observation, we replace direct value regression with a IQL-style target that fits $V$ to the $\tau$-expectile of the critic $Q_{\phi}$:
$\mathcal{L}_{v}(\varphi)=\mathbb{E}_{\bm{s}\sim\mathcal{D}}\!\Bigl{[}L_{2}^{\tau}\!\bigl{(}Q_{\phi}(\bm{s},\bm{a})-V_{\varphi}(\bm{s})\bigr{)}\Bigr{]}.$ (16)
This formulation promotes stable learning by anchoring $V$ to $Q_{\phi}$ and leveraging the advantage-modulated Q-function to improve value estimation. The full procedure is summarized in Algorithm 2. Empirically, this variant achieves value estimates comparable to the original design and demonstrates substantial robustness to the choice of $\tau$ on typical suboptimal offline datasets.
C.6 Computational Efficiency of Training and Inference
Figure 9: Speedup comparison of ADAC, EDP and DQL baseline methods.
All experiments were conducted using NVIDIA RTX 4090 GPUs. Our implementation builds on jaxrl_m, a modular and extensible offline RL codebase based on Flax and JAX. We extend this framework to the proposed Advantage-Based Diffusion Actor-Critic (ADAC).
The optimized ADAC method achieves a training throughput of approximately 115 steps per second, resulting in a 2×2$\times$2 × speedup during training and a 1.64×1.64$\times$1.64 × speedup in inference compared to the original DQL implementation, as shown in Figure 9. ADAC matches the training efficiency of Efficient Diffusion Policy (EDP) [17] while offering superior inference speed. These results highlight the computational efficiency of our approach in both training and inference.
C.7 Auxiliary Model Pretraining
To facilitate offline reinforcement learning, we pretrain two models: a transition model that predicts the next state from a state-action pair, and a behavior cloning model based on a diffusion probabilistic model (DDPM) that learns the action distribution conditioned on the current state. The transition model and the DDPM’s noise predictor are both implemented as MLPs with 256 neurons per layer, using the Mish activation. As shown in Table 4, we use a 95%/5% split, batch size 256, and 300 000 gradient steps, optimized with weight-decayed Adam ($3\times 10^{-4}$ learning rate). Each model trains in under 10 minutes on an NVIDIA RTX 4090. Once trained on a dataset, these models can be reused across experiments, improving efficiency and ensuring consistency regardless of main training hyperparameters.
Table 4: Hyperparameters for Transition and Behavior Cloning Model Pretraining
Hyperparameter Transition Model Behavior Cloning Model
Architecture 4-layer MLP 256 neurons per layer 5-layer MLP 256 neurons per layer
Activation Function Mish Mish
Optimizer AdamW AdamW
Learning Rate $3\times 10^{-4}$ $3\times 10^{-4}$
Batch Size 256 256
Gradient Descent Steps 300 000 300 000
Train/Test Split 95% / 5% 95% / 5%
C.8 Dataset Construction for Offline RL in PointMaze
We utilize the PointMaze environment from Gymnasium Robotics, a refactored version of D4RL’s Maze2D. In this environment, an agent navigates a closed maze using 2D continuous force control (bounded $(x,y)$ forces applied at 10 Hz). For our specific experiments focusing on sparse reward spatial navigation, we simplify the state space by omitting goal-related fields. This results in a compact 4-dimensional observation vector comprising only the agent’s position and velocity.
To construct the offline dataset, we collect trajectories across multiple runs of online SAC [12] training. We employed a staged sampling strategy during online training (Algorithm 3). This strategy involved periodically performing trajectory rollouts using the current policy, allowing us to collect a diverse set of behaviors as the policy evolved throughout training. This collection process spanned 10 independent runs of online SAC training, each for up to 250 000 steps.
Algorithm 3 Trajectory Sampling for Offline Dataset Construction
1:  Initialize: policy $\pi$, environment $\mathcal{E}$, replay buffer $\mathcal{B}$
2:  Define: a staged sampling schedule alternating coarse- and fine-grained rollouts
3:  for each training step do
4:     Interact with $\mathcal{E}$ using $\pi$ and store transitions $(\bm{s_{t}},\bm{a_{t}},r_{t},\bm{s_{t+1}})$ into $\mathcal{B}$
5:     Periodically update $\pi$ using mini-batches sampled from $\mathcal{B}$
6:     if sampling interval is triggered then
7:        Execute coarse-grained or fine-grained trajectory rollout according to schedule
8:        Store resulting trajectories in $\mathcal{D}$
9:     end if
10:  end for
11:  Filter suboptimal trajectories based on return and length heuristics
12:  Return: offline dataset $\mathcal{D}_{\text{offline}}$
In total, the collected offline dataset comprises 853 trajectories. As detailed in Table 5, the dataset exhibits a distribution heavily skewed towards suboptimal behavior, consistent with our collection strategy. Specifically, it contains no optimal trajectories, a very small number of near-optimal paths (3.4%), and a modest proportion (15.7%) of competitive paths, which we consider to be acceptably successful for the task. The vast majority of the dataset (80.9%) consists of trajectories categorized as Sub-Optimal based on their length, defined as exceeding 300 steps.
Table 5: Distribution of trajectory quality based on length in the collected offline dataset.
Category Length Range Proportion
Optimal $<175$ 0.0%
Near-Optimal $[175,225)$ 3.4%
Competitive $[225,300)$ 15.7%
Sub-Optimal $\geq 300$ 80.9%